OpenStudy (anonymous):

How many solutions does the equation y + 12 = y + 10 + 2 have? Zero One Two Infinitely many

pooja195 (pooja195):

y+12=y+10+2 y+12=y+10+2 y+12=(y)+(10+2)(Combine Like Terms) y+12=y+12 y+12=y+12 S Subtract y from both sides. y+12−y=y+12−y 12=12 Subtract 12 from both sides. 12−12=12−12 0=0 when we have a #=# (same number) it is infinte solutions

pooja195 (pooja195):

m=n <---no sol m=m <--infinte m=# <---one